# Statistical Advantage of Softmax Attention: Insights from Single-Location Regression

### v1

@@ A.2.3 Stability of the manifold
starting from a perturbed random initial condition
$(m_{kk^{*}},m_{kv^{*}},m_{vk^{*}},m_{vv^{*}},R_{kk},R_{kv},R_{vv})=(0,0,0,0,1,0,1)+\eta\ ,$
with $\eta\sim\operatorname{Unif}([-\bar{\eta},+\bar{\eta}]^{7})$ some small noise. We consider the value of $\Theta$ reached after convergence, for several independent realizations of $\eta$, for all the configurations considered in the main part of the article. On Figure 3 we observe that all the trajectories converge to a point that belongs to $\tilde{\mathcal{M}}$, up to the numerical errors due to the integration. We additionally observe that $R_{kk}\approx R_{vv}\approx 0$.
@@ A.2.6 Other possible minima
We perform a local minimization of $\tilde{\mathcal{E}}_{\sigma}$ starting from the initial condition
$(m_{kk^{*}},m_{kv^{*}},m_{vk^{*}},m_{vv^{*}},R_{kk},R_{kv},R_{vv})=(0,1,1,0,0,0,0).$
The intuition is that the attention can confuse and mix $k^{*}$ and $v^{*}$. For instance for an identity activation function $\sigma(\chi)=\chi$, $k$ and $v$ play symmetric roles, that we partially broke by adding a constant bias. We call mismatched minimum such a minimum where $m_{kv^{*}}\neq 0,m_{vk^{*}}\neq 0$ and $m_{kk^{*}}=0,m_{vv^{*}}=0$, and set ${\beth}_{\sigma}$ its risk, by opposition with the well matched minimum on $\mathcal{M}$ where $m_{kk^{*}}\neq 0,m_{vv^{*}}\neq 0$ and $m_{kv^{*}}=0,m_{vk^{*}}=0$. On Figure 4 we show that at large enough sequence length $L=10$ and at $\nu>0$, on the two models, the attention always has a well matched and a mismatched minimum, and that the mismatched minimum has a larger population error ${\beth}_{\sigma}>\mathsf{E}_{\sigma}$.
@@ C.2 Replica, free entropy
and similarly for the conjugates. The set of order parameters becomes
$\Theta=(m_{k},\hat{m}_{k},m_{v},\hat{m}_{v},q_{k},\hat{q}_{k},q_{v},\hat{q}_{v},Q_{k},\hat{Q}_{k},Q_{v},\hat{Q}_{v})\in\mathbb{R}^{12}.$
The free entropy is now
@@ D.1 Training of the attention
For most of the parameters of the data model we explored the training converges to the global minimum of the training loss eq. (15), in the sense that the training loss of trained attention is equal to the training loss predicted by our asymptotic characterization.
@@ Appendix E Additional figures
Figure 6: Test error of the trained attention, linear (top) and softmax (bottom), across different tasks and signal strengths $\nu$, for $L=3$. Solid lines indicate $\mathsf{E}_{\sigma}(\alpha)$ at finite $\alpha$ (Result 5.1), while markers represent the test error of an ERM approximated via a local optimization method with $\sqrt{ND}=10^{4}$ and averaged over ten instances. Random initialization means that the attention is initialized at random $k$ and $v$ while informed initialization means it is initialized at $k=k^{*},v=v^{*}$. Dashed lines correspond to the value of $\mathsf{E}_{\sigma}$ in the infinite-$\alpha$ limit (see closed-formed formulas in Proposition 4.1 for softmax and Appendix A.2.7 for linear). The Bayes-optimal risk is shown in black (see Appendix B for a discussion on its discontinuity). Appendix D includes more experimental details.

### v2

@@ A.2.3 Stability of the manifold
starting from a perturbed random initial condition $(m_{kk^{*}},m_{kv^{*}},m_{vk^{*}},m_{vv^{*}},R_{kk},R_{kv},R_{vv})=(0,0,0,0,1,0,1)+\eta$, with $\eta\sim\operatorname{Unif}([-\bar{\eta},+\bar{\eta}]^{7})$ some small noise. We consider the value of $\Theta$ reached after convergence, for several independent realizations of $\eta$, for all the configurations considered in the main part of the article. On Figure 4 we observe that all the trajectories converge to a point that belongs to $\tilde{\mathcal{M}}$, up to the numerical errors due to the integration. We additionally observe that $R_{kk}\approx R_{vv}\approx 0$.
@@ A.2.6 Other possible minima
We perform a local minimization of $\tilde{\mathcal{E}}_{\sigma}$ starting from the initial condition $(m_{kk^{*}},m_{kv^{*}},m_{vk^{*}},m_{vv^{*}},R_{kk},R_{kv},R_{vv})=(0,1,1,0,0,0,0)$. The intuition is that the attention can confuse and mix $k^{*}$ and $v^{*}$. For instance for an identity activation function $\sigma(\chi)=\chi$, $k$ and $v$ play symmetric roles, that we partially broke by adding a constant bias. We call mismatched minimum such a minimum where $m_{kv^{*}}\neq 0,m_{vk^{*}}\neq 0$ and $m_{kk^{*}}=0,m_{vv^{*}}=0$, and set ${\beth}_{\sigma}$ its risk, by opposition with the well matched minimum on $\mathcal{M}$ where $m_{kk^{*}}\neq 0,m_{vv^{*}}\neq 0$ and $m_{kv^{*}}=0,m_{vk^{*}}=0$. On Figure 5 we show that at large enough sequence length $L=10$ and at $\nu>0$, on the two models, the attention always has a well matched and a mismatched minimum, and that the mismatched minimum has a larger population error ${\beth}_{\sigma}>\mathsf{E}_{\sigma}$.
@@ C.2 Replica, free entropy
and similarly for the conjugates. The set of order parameters becomes $\Theta=(m_{k},\hat{m}_{k},m_{v},\hat{m}_{v},q_{k},\hat{q}_{k},q_{v},\hat{q}_{v},Q_{k},\hat{Q}_{k},Q_{v},\hat{Q}_{v})\in\mathbb{R}^{12}$. The free entropy is now
@@ D.1 Training of the attention
For most of the parameters of the data model we explored the training converges to the global minimum of the training loss eq. (15), in the sense that the test risk of trained attention is equal to the test risk predicted by our asymptotic characterization.
@@ Appendix E Additional figures
Figure 7: Minimal test risk $\mathsf{E}_{\sigma}(\alpha)$ of the attention, linear (top) and softmax (bottom), across different tasks and signal strengths $\nu$, for $L=3$. Solid lines indicate $\mathsf{E}_{\sigma}(\alpha)$ (Result 5.1), while markers represent the test risk of an ERM approximated via a local optimization method with $\sqrt{ND}=10^{4}$ and averaged over ten instances. Random initialization means that the attention is initialized at random $k$ and $v$ while informed initialization means it is initialized at $k=k^{*},v=v^{*}$. Dashed lines correspond to the value of $\mathsf{E}_{\sigma}$ in the infinite-$\alpha$ limit (see closed-formed formulas in Proposition 4.1 for softmax and Appendix A.2.7 for linear). The Bayes-optimal risk $\mathcal{E}_{\mathrm{BO}}(\alpha)$ is shown in black (see Appendix B for a discussion on its discontinuity). Appendix D includes more experimental details.